OpenStudy (anonymous):

OKAY!! THIS ONE REALLY CONFUSED ME!!!... what's the common ratio for 2,-6,18,-54?!?!?

OpenStudy (mathstudent55):

Divide one term by the previous term.

OpenStudy (psymon):

-6/2 = 18/-6 = -54/18 = Youll see the pattern :P

OpenStudy (anonymous):

ohh so basically dividing each term with each other? lol with the next!

OpenStudy (psymon):

Pretty much xD

OpenStudy (anonymous):

wow! now why cant they tell you that when your solving it? lol gezz been stuck on this one for quite some time! hahah thank you((:

OpenStudy (mathstudent55):

A geometric series has the same ratio between any to consecutive terms. To find the common ratio, all you need to do is pick any two consecutive terms and divide the later one by the earlier one.